### v3

# Breaking the Low-Rank Dilemma of Linear Attention

Abstract
The Softmax attention mechanism in Transformer models is notoriously computationally expensive due to its quadratic complexity, posing significant challenges in vision applications. In contrast, linear attention offers a far more efficient solution by reducing the complexity to linear levels. However, linear attention often suffers significant performance degradation compared to Softmax attention. Our experiments indicate that this performance drop stems from the low-rank nature of linear attention’s output feature map, which hinders its ability to adequately model complex spatial information. To address this low-rank dilemma, we conduct rank analysis from two perspectives: the KV buffer and the output features. Consequently, we introduce Rank-Augmented Linear Attention (RALA), which rivals the performance of Softmax attention while maintaining linear complexity and high efficiency. Building upon RALA, we construct the Rank-Augmented Vision Linear Transformer (RAVLT). Extensive experiments demonstrate that RAVLT achieves excellent performance across various vision tasks. Specifically, without using any additional labels, data, or supervision during training, RAVLT achieves an 84.4% Top-1 accuracy on ImageNet-1k with only 26M parameters and 4.6G FLOPs. This result significantly surpasses previous linear attention mechanisms, fully illustrating the potential of RALA. Code will be available at https://github.com/qhfan/RALA.
1 Introduction
Figure 1: Comparison of Softmax attention and linear attention. Linear attention has linear complexity and high efficiency, but its spatial modeling capability is inferior to Softmax attention.
In recent years, the Transformer [57] has garnered increasing attention as a powerful foundational architecture. In the field of computer vision, Vision Transformers [13] have made significant breakthroughs in image classification, object detection, instance segmentation, and semantic segmentation, further demonstrating the vast potential of Transformers.
Figure 2: Comparison of feature maps output by Softmax attention and different linear attentions. All experiments are conducted based on the DeiT-T architecture, with $N=196$ and $d=64$. The full rank of matrices in the fig is 64. Compared to Softmax attention, the output features of various linear attentions exhibit significantly low-rank properties. This indicates that the diversity of features learned by linear attention is inferior to that learned by Softmax attention.
Figure 3: Comparison among models based on linear attention and Softmax attention. Our RAVLT achieves state-of-the-art results across all scales and significantly outperforms existing vision models based on linear attention.
However, applying Transformers to vision tasks is not straightforward. Self-attention, as the core mechanism of Transformers, has quadratic complexity. As the number of tokens increases, the computational load of self-attention grows significantly, making Transformers challenging to apply to vision tasks. Many works attempt to address this issue, with some approaches grouping vision tokens to limit the number of tokens each token can interact with [45, 12, 74, 44, 17, 60, 56]. While this method effectively reduces the model’s complexity, it sacrifices the Transformer’s crucial ability to perceive information globally. Another line of work attempts to downsample the keys and values in Softmax attention to capture global features with lower complexity [58, 59, 16, 28]. However, this approach sacrifices the model’s fine-grained perception capabilities.
Unlike the methods mentioned above, linear attention [25, 26, 47, 52, 34, 2] takes a different approach by replacing the Softmax with kernel functions and altering the computation order of $Q$, $K$, and $V$, which is shown in the Fig. 1. This reduces the quadratic complexity of Softmax attention to linear complexity. Additionally, because its computation closely resembles Softmax attention and only uses kernel functions to approximate Softmax, linear attention possesses excellent global modeling capability. Furthermore, since linear attention does not involve downsampling, it retains the model’s fine-grained perception ability.
Despite having a form and advantages very similar to Softmax attention, the actual performance of linear attention is less than satisfactory. Although it is highly efficient, its performance shows a significant gap compared to Softmax attention. To identify the cause of this performance gap, we conduct a comparison based on the DeiT-T architecture, merely replacing different attention mechanisms (vanilla linear attention, Enhanced Linear Attention [2], Efficient Attention [52], Focused Linear Attention [25]), with results shown in Fig. 2. We perform rank analysis on the features of single heads from different output feature maps ($N=196$, $d=64$, with the full rank of the matrix being 64). We find that the rank of output features from existing linear attentions are significantly lower than those from Softmax attention, indicating that the diversity of features they learn is comparatively poorer.
Based on the above considerations, we aim to increase the rank of linear attention to achieve a trade-off between performance and efficiency in vision models. Specifically, we find that two computational steps in linear attention—the calculation of the KV buffer and the calculation of the output features—both affect its low-rank characteristics. Based on this, we propose Rank-Augmented Linear Attention (RALA). In RALA, to address the low-rank issue of the KV buffer, we use a set of context-aware rearrangement coefficients to restructure the weights of each token in the KV buffer. This approach enhances the richness and diversity of information in the KV buffer, which in turn increases the rank of the matrix. For the low-rank issue of the output features, we introduce a feature interaction strategy in the channel dimension, ensuring the output features reach a full-rank state. RALA effectively enhances the rank of the feature matrices in linear attention, enabling it to model complex spatial features with linear complexity. As shown in the Fig. 2, the output feature matrix of RALA reaches the full rank of 64.
Utilizing RALA, we construct the Rank-Augmented Vision Linear Transformer (RAVLT). We conduct extensive experiments with RAVLT on image classification, object detection, instance segmentation, and semantic segmentation, and RAVLT achieves results comparable to state-of-the-art Vision Transformers. RAVLT also exhibits great trade-off between performance and efficiency. As shown in Fig. 3, without using any additional training data, labels, or supervision, our model achieves an accuracy of 84.4% on ImageNet with only 26M parameters and 4.6G FLOPs. As far as we konw, this surpasses all existing vision models.
2 Related works
Vision Transformer.
Transformer is first employed in the field of NLP [57] and have attracted wide interest in computer vision community [13, 73, 20, 62]. The core operation of Transformer is the vanilla Softmax attention, which faces the challenge of quadratic complexity. Many methods have been proposed to address this challenge [27, 23, 63, 15, 44, 51, 19, 11]. Some methods group tokens to limit the receptive field of each token, thereby reducing the complexity of vanilla Softmax attention to linear [45, 17, 12, 44, 67, 54, 32, 60, 56]. Another approach downsamples the key and value, sacrificing some fine-grained perception while preserving the Transformer’s ability to perceive global information, and reduces the complexity of vanilla Softmax attention to linear [58, 59, 23, 16, 63, 4]. Many methods also try to combine classical CNNs with Transformers, leveraging the excellent local perception capabilities of CNNs with the global perception abilities of Transformers to create a powerful vision backbone [15, 30, 49, 23, 24, 68, 53, 35, 9, 40]. All these models rely on the vanilla Softmax attention, which incurs quadratic complexity.
Linear Attention.
Linear attention uses kernel functions to approximate the Softmax function, allowing it to change the computation order of $Q$, $K$, and $V$ in vanilla Softmax attention, thereby reducing the computational complexity from quadratic to linear. Although the efficiency of linear attention is significantly improved compared to vanilla Softmax attention, its performance still lags considerably behind. Many methods have been proposed to improve this [47, 25, 26, 65, 70, 52]. Specifically, FLatten-Transformer [25] proposes focused function to improve the focus ability of linear attention. MLLA [26] connects the linear attention with Mamba [21]. Most of other methods start by using kernel functions to approximate Softmax or introduce DWConv to increase feature diversity. However, the low-rank issue still exists, as shown in Fig. 2.
3 Method
3.1 Preliminary
Softmax Attention.
Softmax attention is the classic mechanism used in Transformers [57] and employed in ViT and its various variants [13, 45, 12, 33, 35, 19]. Specifically, given the input tokens $X\in\mathbb{R}^{N\times d}$, Softmax attention can generally be represented as:
$Y_{i}=\sum_{j=1}^{N}\frac{{\rm Sim}(Q(X_{i}),K(X_{j}))}{\sum_{m=1}^{N}{\rm Sim}(Q(X_{i}),K(X_{m}))}V(X_{j})$ (1)
where $i,j$ are the index of tokens, $X_{i},Y_{i}\in\mathbb{R}^{1\times d}$. $Q(.)$, $K(.)$ and $V(.)$ are the linear transformations for $X$. For simplicity, we use $Q_{i}$ to refer to $Q(X_{i})$, and similarly for $K(.)$ and $V(.)$. In Softmax attention, it is generally expressed as: ${\rm Sim}(Q_{i},K_{j})={\rm exp}(Q_{i}K_{j}^{T}/\sqrt{d})$. Since the computation of Softmax attention must be performed between all queries and keys, its complexity is $O(N^{2}d)$, exhibiting quadratic growth as the number of tokens increases. Therefore, applying it directly in vision models results in significant computational overhead, especially when the number of tokens is large.
Linear Attention.
Compared to Softmax attention, linear attention uses kernel functions to approximate the ${\rm Sim}(.,.)$, allowing the computation of ${\rm Sim}(Q_{i},K_{j})$ to be decomposed into ${\rm Sim}(Q_{i},K_{j})=\kappa(Q_{i})\kappa(K_{j})^{T}$. Specifically, the computation of linear attention is shown as follows:
$Y_{i}$ $=\sum_{j=1}^{N}\frac{{\rm Sim}(Q_{i},K_{j})}{\sum_{m=1}^{N}{\rm Sim}(Q_{i},K_{m})}V_{j}$ (2)
$=\sum_{j=1}^{N}\frac{\kappa(Q_{i})\kappa(K_{j})^{T}}{\sum_{m=1}^{N}\kappa(Q_{i})\kappa(K_{m})^{T}}V_{j}$
$=\frac{\kappa(Q_{i})(\sum_{j=1}^{N}\kappa(K_{j})^{T}V_{j})}{\kappa(Q_{i})(\sum_{m=1}^{N}\kappa(K_{m})^{T})}$
where $\kappa(.)$ is the kernel function. Compared to the computation order $(QK^{T})V$ in Softmax attention, the computation order $Q(K^{T}V)$ in linear attention reduces the complexity with respect to the number of tokens ($N$) from quadratic $O(N^{2}d)$ to linear $O(Nd^{2})$. However, the reduction in complexity also results in a decline in performance. As shown in Fig. 2, the output features of linear attention have a low rank, indicating that linear attention fails to learn diverse visual features.
Figure 4: Visualization of the rank analysis of the KV buffer for different linear attention mechanisms. The KV buffer $\sum_{j=1}^{N}\kappa(K_{j})^{T}V_{j}\in\mathbb{R}^{64\times 64}$.
3.2 Rank-Augmented Linear Attention
In the general expression of linear attention, since the denominator $\kappa(Q_{i})(\sum_{j=1}^{N}\kappa(K_{j})^{T})\in\mathbb{R}^{1\times 1}$, the rank of the output feature matrix $Y$ only depends on the numerator. Therefore, in the subsequent analysis, we focus solely on the numerator $\kappa(Q_{i})(\sum_{j=1}^{N}\kappa(K_{j})^{T}V_{j})$.
Augment the rank of the KV buffer.
Each $Q_{i}$ needs to be multiplied by the KV buffer ($\sum_{j=1}^{N}\kappa(K_{j})^{T}V_{j}$) to obtain the final $Y_{i}$. Therefore, the diversity of the features in the KV buffer, reflected by its rank, dictates the diversity of the final output features. Thus, we first analyze the rank of the KV buffer. As shown in Fig. 4, the rank of the KV buffer is relatively low in almost all linear attentions. This means that the KV buffer does not store sufficient information for each $Q_{i}$ to query. The feature diversity in the KV buffer is poor.
To increase the rank of the KV buffer, we introduce operations into its calculation to eliminate linear dependencies between different rows and columns. Beyond the feature diversity, the rank of the KV buffer is also closely related to the amount of information it contains, increasing the information content in the KV buffer will simultaneously raise its rank. Based on this consideration, rather than assigning a weight of 1 to each $\kappa(K_{j})^{T}V_{j},j\in[1,N]$ as in other linear attentions [52, 25], we aim to highlight the significance of each token in the KV buffer by assigning higher weights to the tokens that contain more information. That is:
$B=\sum_{j=1}^{N}\alpha_{j}\kappa(K_{j})^{T}V_{j},\quad\sum_{j=1}^{N}\alpha_{j}=N$ (3)
where $B$ denotes the KV buffer, $\alpha_{j}$ is the weight coefficient of the j-th token.
To determine the value of $\alpha_{j}$, we draw inspiration from the design of Softmax attention. In Softmax attention, the attention scores between tokens are generally considered to represent the degree of correlation between them. Tokens more correlated with the global token are considered to contain more important information[41]. Based on this fact, by calculating the attention score of the global query to each key, we determine the proportion of each token in the KV buffer, resulting in the following:
$Q_{g}=\frac{\sum_{i=1}^{N}Q_{i}}{N}$ (4)
$\alpha_{j}$ $=N\times\frac{{\rm exp}(Q_{g}\kappa(K_{j})^{T})}{\sum_{m=1}^{N}{\rm exp}(Q_{g}\kappa(K_{m})^{T})}$
As shown in Fig. 4, after introducing $\alpha_{j}$ as the modulation coefficient for $\kappa(K_{j})^{T}V_{j}$, the information content in the KV buffer increases, which significantly enhances its rank. More analysis about why the introduction of $\alpha_{j}$ can increase the matrix’s rank can be found in Appendix.
Augment the rank of the output features.
In the fundamentals of matrix theory, for matrices $C$ and $D$, we have: ${\rm Rank}(CD)\leq{\rm min}({\rm Rank}(C),\>{\rm Rank}(D))$. Therefore, when calculating the output features $Y=\kappa(Q)B$, its rank:
${\rm Rank}(Y)\leq{\rm min}({\rm Rank}(\kappa(Q)),\>{\rm Rank}(B))\leq d$ (5)
$Y\in$ $\mathbb{R}^{N\times d},\>\kappa(Q)\in\mathbb{R}^{N\times d},\>B=\sum_{j=1}^{N}\alpha_{j}\kappa(K_{j})^{T}V_{j}\in\mathbb{R}^{d\times d}$
This indicates that even though the KV buffer $B$ reaches a full-rank state, the output feature matrix is still highly likely to be in a low-rank state (${\rm Rank}(Y)<d$), meaning information loss occurs when computing each $Y_{i}$. As shown in Fig. 5, we confirm this point when performing rank analysis on the output features. The output features exhibit a certain degree of rank reduction.
Figure 5: Visualization of the output features’ ($Y\in\mathbb{R}^{N\times d}$, $N=196$, $d=64$) rank analysis.
To restore the output feature matrix $Y$ to a full-rank state, we aim to complete the information loss for each $Y_{i}$.
Figure 6: RAVLT’s block (left) and RALA (right). We employ serial numbers to represent the sequence of calculations. $\odot$ represents Hadamard product, and $\times$ represents matrix multiplication.
Specifically, we post-process each output feature $Y_{i}$ by modulating $Y_{i}$ with a transformed form that contains the information of the original token $X_{i}$. At this stage, $Y_{i}$ has already completed token mixing, so this modulation is applied only along the channel dimension. Consequently, each output feature becomes:
$Y_{i}=\phi(X_{i})\odot(\kappa(Q_{i})\sum_{j=1}^{N}\alpha_{j}\kappa(K_{j})^{T}V_{j})$ (6)
where $\odot$ denotes the Hadamard product. As shown in the Fig. 5, After introducing the modulation coefficients, the feature diversity of the output feature matrix is significantly enhanced, transforming it from a low-rank state to a full-rank state. This indicates that token-specific modulation adds a degree of feature diversity to the output features. More analysis can be found in the Appendix.
Design of the block.
Fig. 6 illustrates the design of the entire block. Our block design follows the classic Transformer block design [57], consisting of linear attention and FFN. All the networks in this paper are constructed by stacking this basic block. We use a simple linear projection to serve as the transformation function $\phi(.)$. As for $\kappa(.)$, we have $\kappa(.)={\rm Elu}(.)+1$. For each block, we use CPE [6] to serve as the positional encoding. The CPE is a simple $3\times 3$ depth-wise convolution with residual connection.
3.3 Implementation Details
Model Blocks Channels Heads Params (M) FLOPs (G)
RAVLT-T [2,2,6,2] [64,128,256,512] [1,2,4,8] 15 2.4
RAVLT-S [3,5,9,3] [64,128,320,512] [1,2,5,8] 26 4.6
RAVLT-B [4,6,12,6] [96,192,384,512] [1,2,6,8] 48 9.9
RAVLT-L [4,7,19,8] [96,192,448,640] [1,2,7,10] 95 16.0
Table 1: Architecture variants of RAVLT. The FLOPs are measured at resolution $224\times 224$.
Based on the block in the Fig. 6, we construct the Rank-Augmented Vision Linear Transformer (RAVLT), a hierarchical general vision backbone. As shown in the Tab. 1, following the previous works [19, 74, 33], we build three RAVLT backbones with different settings of block number and channel number in each stage. In each stage, the image downsampling rates are $\frac{1}{4}$, $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$, respectively. At the beginning of each stage, we use a $3\times 3$ convolution with the stride of $2$ to downsample the image.
4 Experiments
We conduct extensive experiments on multiple vision tasks, such as image classification, object detection, instance segmentation, and semantic segmentation. We also make ablation studies to validate the importance of each components in RALA. More details can be found in Appendix.
4.1 Image Classification
Settings.
We train our models on ImageNet-1K [10] from scratch and follow the same training strategy in DeiT [55] with the only supervision being classification loss. The maximum rates of increasing stochastic depth [31] are set to 0.1/0.15/0.4/0.55 for RAVLT-T/S/B/L, respectively. We evaluate the models on ImageNet [10] and ImageNet-V2 [50].
Figure 7: Comparison of multiple general backbones’ inference speed. The speed are measured on A100. RAVLT exhibits the best trade-off between efficiency and performance.
Cost Model Params(M) FLOPs(G) Top1-acc(%)
Tiny model $\sim 2.5$G tiny-MOAT-2 [66] 10 2.3 70.1
BiFormer-T [74] 13 2.2 70.7
SMT-T [43] 12 2.4 71.0
RMT-T [19] 14 2.5 72.1
RAVLT-T 15 2.4 72.7
Small model $\sim 4.5$G DeiT-S 22 4.6 68.5
MOAT-0 [66] 28 5.7 72.8
SMT-S [43] 21 4.7 73.3
MLLA-T [26] 25 4.2 73.3
BiFormer-S [74] 26 4.5 73.6
RMT-S [19] 27 4.5 74.1
RAVLT-S 26 4.6 74.9
Base model $\sim 10.0$G XCiT-S24 [14] 48 9.1 73.3
BiFormer-B [74] 57 9.8 74.0
MOAT-1 [66] 42 9.1 74.2
MLLA-S [26] 43 7.3 74.9
RMT-B [19] 54 9.7 75.6
RAVLT-B 48 9.9 76.3
Large model $\sim 15.0$G DeiT-B [55] 86 17.5 71.5
MOAT-2 [66] 73 17.2 74.3
RMT-L [19] 95 18.2 76.3
MLLA-B [26] 96 16.2 76.7
RAVLT-L 95 16.0 76.8
Table 2: Classification results of different general backbones on ImageNet-V2
Cost Model Type Parmas (M) FLOPs (G) Top1-acc (%)
Tiny model $\sim 2.5$G VAN-B1 [24] CNN 14 2.5 81.1
Conv2Former-N [30] CNN 15 2.2 81.5
PVTv2-b1 [59] Trans 13 2.1 78.7
MPViT-XS [37] Trans 11 2.9 80.9
tiny-MOAT-2 [66] Trans 10 2.3 81.0
NAT-M [27] Trans 20 2.7 81.8
GC-ViT-XT [28] Trans 20 2.6 82.0
SMT-T [43] Trans 12 2.4 82.2
RMT-T [19] Trans 14 2.5 82.4
FL-PVT-T [25] Linear 12 2.0 77.8
SOFT-Tiny [47] Linear 13 1.9 79.3
FL-PVTv2-B1 [25] Linear 13 2.2 79.5
RAVLT-T Linear 15 2.4 82.8
Small model $\sim 4.5$G ConvNeXT-T [46] CNN 29 4.5 82.1
FocalNet-T [68] CNN 29 4.5 82.3
MogaNet-S [39] CNN 25 5.0 83.4
InternImage-T [61] CNN 30 5.0 83.5
Swin-T [45] Trans 29 4.5 81.3
SG-Former-S [22] Trans 23 4.8 83.2
StructViT-S-8-1 [35] Trans 24 5.4 83.3
MOAT-0 [66] Trans 28 5.7 83.3
FAT-B3 [16] Trans 29 4.4 83.6
SMT-S [43] Trans 20 4.8 83.7
BiFormer-S [74] Trans 26 4.5 83.8
RMT-S [19] Trans 27 4.5 84.1
FL-Swin-T [25] Linear 29 4.5 82.1
SOFT-Small [47] Linear 24 3.3 82.2
FL-CSwin-T [25] Linear 21 4.3 83.1
MLLA-T [26] Linear 25 4.2 83.5
RAVLT-S Linear 26 4.6 84.4
Cost Model Type Parmas (M) FLOPs (G) Top1-acc (%)
Base model $\sim 10.0$G ConvNeXT-S [46] CNN 50 8.7 83.1
InternImage-S [61] CNN 50 8.0 84.2
MogaNet-B [39] CNN 44 9.9 84.3
ScaleViT-B [69] Trans 81 8.6 84.1
DaViT-S [11] Trans 50 8.8 84.2
StructViT-B-8-1 [35] Trans 52 12.0 84.3
BiFormer-B [74] Trans 57 9.8 84.3
GC-ViT-S [28] Trans 51 8.5 84.3
STViT-B [33] Trans 52 9.9 84.8
RMT-B [19] Trans 54 9.7 85.0
SOFT-Large [47] Linear 64 11.0 83.1
FL-Swin-S [25] Linear 51 8.7 83.5
FL-CSwin-S [25] Linear 35 6.9 83.8
MLLA-S [26] Linear 43 7.3 84.4
RAVLT-B Linear 48 9.9 85.5
Large model $\sim 15.0$G HorNet-B [49] CNN 88 15.5 84.3
MogaNet-L [39] CNN 83 15.9 84.7
InterImage-B [61] CNN 97 16.0 84.9
NAT-B [27] Trans 90 13.7 84.3
SMT-L [43] Trans 81 17.7 84.6
SG-Former-B [22] Trans 78 15.6 84.7
GC-ViT-B [28] Trans 90 14.8 85.0
FAT-B5 [16] Trans 88 15.1 85.2
STViT-L [33] Trans 95 15.6 85.3
RMT-L [19] Trans 95 18.2 85.5
SOFT-Huge [47] Linear 87 16.3 83.3
FLatten-Swin-B [25] Linear 89 15.4 83.8
FLatten-CSwin-B [25] Linear 75 15.0 84.5
MLLA-B [26] Linear 96 16.2 85.3
RAVLT-L Linear 95 16.0 85.8
Table 3: Comparison with the state-of-the-art on ImageNet-1K classification. We use ”CNN” to refer to convolutional neural networks, ”Trans” to refer to Vision Transformers, and ”Linear” to refer to models based on linear operators.
Results.
In Tab. 3 and Tab. 2, we compare the performance of various models on ImageNet and ImageNet-v2, respectively. Under models of comparable size, RAVLT achieves the best results. Specifically, RAVLT-B surpasses RMT-B by 0.5% with the same FLOPs. This result demonstrates the superior of RAVLT.
4.2 Inference Speed
We evaluate the RAVLT’s inference efficiency on A100, the results are shown in the Fig. 7. It can be seen that our models exhibit better trade-off between performance and efficiency than other models.
4.3 Object Detection and Instance Segmentation
Settings.
We adopt MMDetection [5] to implement RetinaNet [42], Mask-RCNN [29] and Cascade Mask R-CNN [3]. Following previous works [19, 45, 58], we use the commonly used “1×1$\times$1 ×” (12 training epochs) setting for the RetinaNet and Mask R-CNN. Besides, we use “$3\times+\mathrm{MS}$” for Mask R-CNN and Cascade Mask R-CNN.
Backbone Type Params (M) FLOPs (G) Mask R-CNN 3×3$\times$3 ×+MS
$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
ConvNeXt-T [46] CNN 48 262 46.2 67.9 50.8 41.7 65.0 45.0
NAT-T [27] Trans 48 258 47.8 69.0 52.6 42.6 66.0 45.9
GC-ViT-T [28] Trans 48 291 47.9 70.1 52.8 43.2 67.0 46.7
SMT-S [43] Trans 40 265 49.0 70.1 53.4 43.4 67.3 46.7
RMT-S [19] Trans 46 262 50.7 71.9 55.6 44.9 69.1 48.4
FL-Swin-T [25] Linear 49 268 46.5 68.5 50.8 42.1 65.4 45.1
MLLA-T [26] Linear 44 255 48.8 71.0 53.6 43.8 68.0 46.8
RAVLT-S Linear 44 262 51.4 72.3 56.5 45.5 69.7 48.8
ConvNeXt-S [46] CNN 70 348 47.9 70.0 52.7 42.9 66.9 46.2
InternImage-S [61] CNN 69 340 49.7 71.1 54.5 44.5 68.5 47.8
NAT-S [27] Trans 70 330 48.4 69.8 53.2 43.2 66.9 46.4
Swin-S [45] Trans 69 359 48.5 70.2 53.5 43.3 67.3 46.6
SMT-B [43] Trans 52 328 49.8 71.0 54.4 44.0 68.0 47.3
RMT-B [19] Trans 73 373 52.2 72.9 57.0 46.1 70.4 49.9
MLLA-S [26] Linear 63 319 50.5 71.8 55.2 44.9 69.1 48.2
RAVLT-B Linear 67 372 52.7 73.5 57.7 46.4 70.6 50.2
InternImage-B [61] CNN 115 501 50.3 71.4 55.3 44.8 68.7 48.0
Swin-B [45] Trans 107 496 48.6 70.0 53.4 43.3 67.1 46.7
CSwin-B [12] Trans 97 526 50.8 72.1 55.8 44.9 69.1 48.3
RAVLT-L Linear 114 501 53.6 74.4 58.9 47.3 71.6 51.2
Table 4: Comparison to other backbones using Mask R-CNN with ”$3\times+\mathrm{MS}$” schedule.
Backbone Type Params (M) FLOPs (G) Cascade Mask R-CNN 3×3$\times$3 ×+MS
$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
HorNet-T [49] CNN 80 728 52.4 71.6 56.8 45.6 69.1 49.6
GC-ViT-T [28] Trans 85 770 51.6 70.4 56.1 44.6 67.8 48.3
SMT-S [43] Trans 78 744 51.9 70.5 56.3 44.7 67.8 48.6
UniFormer-S [38] Trans 79 747 52.1 71.1 56.6 45.2 68.3 48.9
CSWin-T [12] Trans 80 757 52.5 71.5 57.1 45.3 68.8 48.9
RMT-S [19] Trans 83 741 53.2 72.0 57.8 46.1 69.8 49.8
FL-Swin-T [25] Linear 87 747 50.8 69.6 55.1 44.1 67.0 48.1
RAVLT-S Linear 82 741 54.2 72.9 58.7 46.8 70.5 50.9
HorNet-S [49] CNN 108 827 53.3 72.3 57.8 46.3 69.9 50.4
NAT-S [27] Trans 108 809 51.9 70.4 56.2 44.9 68.2 48.6
GC-ViT-S [28] Trans 108 866 52.4 71.0 57.1 45.4 68.5 49.3
DAT-S [63] Trans 107 857 52.7 71.7 57.2 45.5 69.1 49.3
UniFormer-B [38] Trans 107 878 53.8 72.8 58.5 46.4 69.9 50.4
RMT-B [19] Trans 111 852 54.5 72.8 59.0 47.2 70.5 51.4
FL-Swin-S [25] Linear 108 841 52.2 71.2 56.8 45.4 68.3 49.4
RAVLT-B Linear 105 851 55.3 73.8 60.1 47.7 71.4 52.1
ConvNeXt-B [46] CNN 145 964 52.7 71.3 57.2 45.6 68.9 49.5
Swin-B [45] Trans 145 982 51.9 70.9 56.5 45.0 68.4 48.7
CSwin-B [12] Trans 135 1004 53.9 72.6 58.5 46.4 70.0 50.4
RAVLT-L Linear 152 979 55.6 74.1 60.5 48.0 71.8 52.3
Table 5: Comparison to other backbones using Cascade Mask R-CNN with ”$3\times+\mathrm{MS}$” schedule.
Backbone Type Params (M) FLOPs (G) Mask R-CNN 1×1$\times$1 × Params (M) FLOPs (G) RetinaNet 1×1$\times$1 ×
$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$ $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{b}_{S}$ $AP^{b}_{M}$ $AP^{b}_{L}$
PVTv2-B1 [59] Trans 33 243 41.8 54.3 45.9 38.8 61.2 41.6 23 225 41.2 61.9 43.9 25.4 44.5 54.3
MPViT-XS [37] Trans 30 231 44.2 66.7 48.4 40.4 63.4 43.4 20 211 43.8 65.0 47.1 28.1 47.6 56.5
SOFT++-Tiny [48] Linear 32 212 41.2 63.7 44.7 38.2 61.0 41.0 23 200 41.9 62.7 44.7 27.8 45.4 55.6
FL-PVT-T [25] Linear 32 244 38.2 61.6 41.9 37.0 57.6 39.0
RAVLT-T Linear 33 219 47.3 69.1 51.9 42.7 66.2 46.0 24 201 45.9 67.4 49.4 28.5 50.1 60.8
InternImage-T [61] CNN 49 270 47.2 69.0 52.1 42.5 66.1 45.8
CMT-S [23] Trans 45 249 44.6 66.8 48.9 40.7 63.9 43.4 44 231 44.3 65.5 47.5 27.1 48.3 59.1
RMT-S [19] Trans 46 262 49.0 70.8 53.9 43.9 67.8 47.4 36 244 47.8 69.1 51.8 32.1 51.8 63.5
FL-Swin-T [25] Linear 49 268 44.2 67.3 48.5 40.2 63.8 43.0
MLLA-T [26] Linear 44 255 46.8 69.5 51.5 42.1 66.4 45.0
RAVLT-S Linear 44 262 49.8 71.3 54.5 44.6 68.5 48.2 34 244 48.3 69.8 52.1 32.7 52.8 63.6
InternImage-S [61] CNN 69 340 47.8 69.8 52.8 43.3 67.1 46.7
CSWin-S [12] Trans 54 342 47.9 70.1 52.6 43.2 67.1 46.2
STViT-B [33] Trans 70 359 49.7 71.7 54.7 44.8 68.9 48.7
SOFT++-medium [48] Linear 69 342 46.6 67.8 51.2 42.0 64.8 45.2 59 322 44.3 64.7 47.4 29.0 48.2 59.9
MLLA-S [26] Linear 63 319 49.2 71.5 53.9 44.2 68.5 47.2
RAVLT-B Linear 67 372 51.2 72.7 56.4 45.7 69.9 49.5 57 353 49.8 71.2 54.0 34.0 54.3 64.9
InternImage-B [61] CNN 115 501 48.8 70.9 54.0 44.0 67.8 47.4
Swin-B [45] Trans 107 496 46.9 69.2 51.6 42.3 66.0 45.5 98 477 45.0 66.4 48.3 28.4 49.1 60.6
Focal-B [67] Trans 110 533 47.8 70.2 52.5 43.2 67.3 46.5 101 514 46.3 68.0 49.8 31.7 50.4 60.8
MPViT-B [37] Trans 95 503 48.2 70.0 52.9 43.5 67.1 46.8 85 482 47.0 68.4 50.8 29.4 51.3 61.5
RMT-L [19] Trans 114 557 51.6 73.1 56.5 45.9 70.3 49.8 104 537 49.4 70.6 53.1 34.2 53.9 65.2
MLLA-B [26] Linear 115 502 50.5 72.0 55.4 45.0 69.3 48.6
RAVLT-L Linear 114 501 52.3 73.8 57.3 46.4 71.1 50.4 104 482 50.9 72.2 55.0 34.7 55.7 65.4
Table 6: Comparison to other backbones using RetinaNet and Mask R-CNN with “1×1$\times$1 ×” schedule.
Results.
[…]
4.4 Semantic Segmentation
Settings.
We adopt the Semantic FPN [36] and UperNet [64] based on MMSegmentation [7], apply RAVLTs which are pretrained on ImageNet-1K as backbone. We use the same setting of PVT [58] to train the Semantic FPN, and follow the default settings in Swin [45] to train the UperNet.
Model Type Semantic FPN 80K Upernet 160K
Params (M) FLOPs (G) mIoU (%) Params (M) FLOPs (G) mIoUss (%)
VAN-B1 [24] CNN 18 140 42.9
PVTv2-B1 [59] Trans 18 136 42.5
RMT-T [19] Trans 17 136 46.4
FL-PVT-T [25] Linear 16 169 37.2
RAVLT-T Linear 18 136 47.9 44 893 49.3
MogaNet-S [39] CNN 29 189 47.7 55 946 49.2
StructViT-S [35] Trans 26 271 46.9
SMT-S [43] Trans 50 935 49.2
SGFormer-S [22] Trans 25 205 49.0 53 989 49.9
RMT-S [19] Trans 30 180 49.4 56 937 49.8
Fl-Swin-T [25] Linear 60 946 44.8
RAVLT-S Linear 28 180 49.5 55 937 50.7
MogaNet-B [39] CNN 74 1050 50.1
InterImage-S [61] CNN 80 1017 50.2
DAT-S [63] Trans 53 320 46.1 81 1079 48.3
StructViT-B [35] Trans 54 529 48.5
RMT-B [19] Trans 57 294 50.4 83 1051 52.0
FL-Swin-S [25] Linear 82 1038 48.1
RAVLT-B Linear 51 292 51.9 77 1050 52.5
MogaNet-L [39] CNN 113 1176 50.9
CSWin-B [12] Trans 81 464 49.9 109 1222 51.1
SGFormer-B [22] Trans 81 475 50.6 109 1304 52.0
RMT-L [19] Trans 98 482 51.4 125 1241 52.8
MLLA-B [26] Linear 128 1183 51.9
RAVLT-L Linear 98 424 52.6 125 1182 53.2
Table 7: Comparison with the state-of-the-art on ADE20K.
Results.
We show the segmentation results in the Tab. 7. RAVLT surpasses its competitors across various frameworks. For example, with the framework of Semantic FPN, RAVLT-B achieves 51.9mIoU. The result even surpasses larger RMT-L. With the framework of UperNet, RAVLT can achieve the mIoU of 53.2.
4.5 Comparison with Other Linear Attention
For a fair comparison with other linear attention mechanisms, similar to FLatten-Transformer [25], we use the DeiT-T [55], Swin-T, and Swin-S [45] network architectures, replacing their attention mechanisms with linear attention mechanisms. As shown in Tab. 8, our RALA significantly outperforms other linear attention mechanisms and also exceeds the performance of Softmax attention.
Linear Attention Params(M) FLOPs(G) Top1-acc(%)
Comparison on DeiT-T Setting
DeiT-T [55] 6 1.1 72.2
Hydra Attn [1] 6 1.1 68.3
Efficient Attn [52] 6 1.1 70.2
Linear Angular Attn [70] 6 1.1 70.8
Enhanced Linear Attn [2] 6 1.1 72.9
Focused Linear Attn [25] 6 1.1 74.1
Ours 6 1.1 75.1
Comparison on Swin-T Setting
Swin-T [45] 29 4.5 81.3
Hydra Attn [1] 29 4.5 80.7
Efficient Attn [52] 29 4.5 81.0
Linear Angular Attn [70] 29 4.5 79.4
Enhanced Linear Attn [2] 29 4.5 81.8
Focused Linear Attn [25] 29 4.5 82.1
Ours 29 4.5 83.4
Comparison on Swin-S Setting
Swin-S [45] 50 8.7 83.0
Focused Linear Attn [25] 51 8.7 83.5
Ours 50 8.7 85.0
Table 8: Comparison of different linear attention based on DeiT-T, Swin-T, and Swin-S. RALA surpasses others by a large margin.
4.6 Ablation Study
In this section, we conduct ablation experiments on various key modules of the model to verify the performance enhancements contributed by each component.
Augmentation of the KV buffer.
We first ablate the effect of augmenting the KV buffer on the model. As shown in Tab. 9, augmenting the KV buffer significantly increases its rank and enhances feature diversity, thereby leading to substantial improvements in model performance. This enhancement is particularly evident in downstream tasks such as object detection.
Augmentation of the output features.
Since we use a linear projection to serve as $\phi(.)$, ablating the augmentation of the output feature reduces the model’s parameter count. To ensure a fair comparison, we increase the expansion ratio of the FFN to maintain the model’s parameter count and FLOPs at the same level. As shown in Tab. 9, the introduction of augmentation effectively increases the rank of the output features, thereby enhancing model’s performance.
Conditional Position Encoding.
We use CPE [6] to provide the positional information to our model. As shown in Tab. 9, CPE also improves the model’s performance to some extent across various tasks. This indicates that CPE provides effective positional information for RAVLT.
Selection of $\phi(.)$ and $\kappa(.)$.
We compare the choices of functions $\phi(.)$ (linear projection, identity map, and tanh) and $\kappa(.)$ (${\rm Elu}(.)+1$ and ${\rm ReLU}(.)$), and experiments show that the model’s performance improves with the inclusion of $\phi(.)$, regardless of the specific choice of it. This conclusion also holds for $\kappa(.)$, as shown in Tab. 9. This indicates that the performance improvements stem from our RALA’s augmentation of matrix’s rank, rather than the selection of specific functions.
Params FLOPs Top1 $AP^{b}$ $AP^{m}$ mIoU
15 2.4 82.8 47.3 42.7 47.9
w/o KV aug 15 2.4 82.1(-0.7) 43.7(-3.6) 39.0(-3.7) 43.6(-4.3)
w/o out aug 15 2.4 82.5(-0.3) 46.3(-1.0) 41.6(-1.1) 47.0(-0.9)
w/o CPE 15 2.4 82.7(-0.1) 47.0(-0.3) 42.5(-0.2) 47.6(-0.3)
$\phi$: $\xrightarrow{}$ identity 15 2.4 82.8(-0.0) 47.2(-0.1) 42.5(-0.2) 48.1(+0.2)
$\phi$: $\xrightarrow{}$ tanh 15 2.4 82.7(-0.1) 47.2(-0.1) 42.7(-0.0) 47.7(-0.2)
$\kappa$: ${\rm ELU}+1$ $\xrightarrow{}{\rm ReLU}$ 15 2.4 82.8(-0.0) 47.3(-0.0) 42.8(+0.1) 47.7(-0.2)
$\kappa$: ${\rm ELU}+1$ $\xrightarrow{}{\rm Softmax}$ 15 2.4 82.7(-0.1) 47.1(-0.2) 42.7(+0.0) 47.8(-0.1)
Table 9: Ablation on RAVLT-T. The results show that performance improvements stem from our RALA’s augmentation of matrix’s rank, rather than the selection of specific functions.
Rank analysis of different layers.
Under the DeiT-T setting, we analyze the rank of different layers of our model (12 layers in total, with each layer’s feature matrix having a shape of $N=196$, $d=64$, and a full rank of 64). As shown in Fig. 8, RALA effectively increases the rank of the model’s output features, thereby enhancing its performance.
Figure 8: Analysis of the rank across different model layers. RALA’s augmentation of the KV buffer and output feature effectively increases the rank of the output features at each layer.
5 Conclusion
In this paper, we investigate the causes of performance degradation in linear attention and conclude that the low rank of its output features is the primary factor. To address this issue, we propose Rank-Augmented Linear Attention (RALA). This method enhances two critical computation steps in linear attention—the calculation of the KV buffer and the computation of the output features—by employing two transformations. These augmentations effectively increase the rank of the matrices, thereby mitigating the low-rank problem in linear attention. Building upon RALA, we develop the Rank-Augmented Vision Linear Transformer (RAVLT), a novel architecture based on linear attention. RAVLT outperforms many existing Transformers, demonstrating significant potential for practical applications.
Appendix A More Analysis About $\alpha$
In this section, we further analyze the rank-boosting effect of $\alpha$ on the KV buffer and conduct additional ablation experiments on $\alpha$ to validate the rationale behind our design.
An analysis of the rank-boosting effect.
Assume in $\kappa(K_{j}^{T})V_{j},j\in[0,N]$ that the two components $\kappa(K_{1}^{T})V_{1}$ and $\kappa(K_{2}^{T})V_{2}$ are linearly correlated. That means:
$\kappa(K_{1}^{T})V_{1}=\lambda\kappa(K_{2}^{T})V_{2}.$ (7)
In general linear attention mechanisms, a portion of the KV buffer formed by the sum of the two can be expressed as:
$\kappa(K_{1}^{T})V_{1}+\kappa(K_{2}^{T})V_{2}=(1+\lambda)\kappa(K_{2}^{T})V_{2}$ (8)
After introducing $\alpha$ as a modulation coefficient for each term, their sum becomes:
$\alpha_{1}\kappa(K_{1}^{T})V_{1}+\alpha_{2}\kappa(K_{2}^{T})V_{2}$ $=\lambda\alpha_{1}\kappa(K_{2}^{T})V_{2}+\alpha_{2}\kappa(K_{2}^{T})V_{2}$ (9)
$=\alpha_{2}(1+\frac{\lambda\alpha_{1}}{\alpha_{2}})\kappa(K_{2}^{T})V_{2}$
Since in our setup the value of $\alpha_{1}$ and $\alpha_{2}$ are calculated based on the attention scores between global query and keys, it changes with the input samples and the training process. This results in more varied coefficients, meaning that the new matrix composed of these two linearly correlated matrices has a broader range of possible values and greater flexibility. This makes the KV buffer $\sum_{j=1}^{N}\alpha_{j}\kappa(K_{j}^{T})V_{j}$ more likely to achieve a full-rank state.
As previously mentioned (Eq. 8 and Eq. 9), two linearly correlated matrices can be summed to form a single matrix with coefficients. Therefore, we consider here only the case where all matrices are linearly independent. Consider the equation:
$\sum_{j=1}^{N}c_{j}\kappa(K_{j}^{T})V_{j}=C$ (10)
Here, $C\in\mathbb{R}^{d\times d}$ is a full-rank matrix. Since all $\kappa(K_{j}^{T})V_{j}$ are linearly independent, this equation has a unique solution or no solution. After decomposing $c_{j}$ into $d_{j}\alpha_{j}$, the original equation becomes:
$\sum_{j=1}^{N}d_{j}\alpha_{j}\kappa(K_{j}^{T})V_{j}=C$ (11)
Due to the presence of different scalars, the solutions can be multiple, giving the matrix representation a broader scope. This implies that there are more solutions that enable the KV buffer to achieve a full-rank state.
More choices about $\alpha$.
We choose the attention scores between the keys and the global query as the value for $\alpha$, and here we experiment with different selections. When $\alpha$ is set as a learnable vector, its fixed shape of $1\times N$ makes it challenging to apply to higher resolution tasks such as object detection. Therefore, we only evaluate its performance on image classification tasks.
Model Params(M) FLOPs(G) Acc(%)
DeiT-T 6 1.1 72.2
attn score 6 1.1 75.1
learnable 6 1.1 73.8(-1.3)
Table 10: More choice about $\alpha$.
We conduct experiments using the DeiT-T [55] configuration, and the results are shown in Tab. 10. While the learnable $\alpha$ still significantly enhances the model’s performance, its effectiveness is not as impressive as the $\alpha$ based on the attention score.
Appendix B More Analysis About $\phi(.)$
An analysis of the rank-boosting effect.
Consider two matrices $A$ and $B$, $A,B\in\mathbb{R}^{m\times n}$, and represent these matrices as a linear combination of rank-one matrices. That is:
$A=\sum_{i=1}^{r}u_{i}v_{i}^{T},\quad B=\sum_{j=1}^{s}x_{j}y_{j}^{T}$ (12)
where $r={\rm Rank}(A)$, $s={\rm Rank}(B)$, $u_{i},x_{j}\in\mathbb{R}^{m\times 1}$, $v_{i},y_{j}\in\mathbb{R}^{n\times 1}$. Consider the Hardmard product:
$(A\odot B)_{ij}=A_{ij}\odot B_{ij}$ (13)
we can rewrite it as:
$A\odot B$ $=(\sum_{i=1}^{r}u_{i}v_{i}^{T})\odot(\sum_{j=1}^{s}x_{j}y_{j}^{T})$ (14)
$=\sum_{i=1}^{r}\sum_{j=1}^{s}(u_{i}\odot x_{j})(v_{i}\odot y_{j})^{T}$
This expression shows that $A\odot B$ can be viewed as a linear combination of $rs$ rank-one matrices. Therefore, the total rank of $A\odot B$ does not exceed the number of matrices, which is:
${\rm Rank}(A\odot B)\leq rs={\rm Rank}(A)\times{\rm Rank}(B)$ (15)
The above expression effectively demonstrates that when the ranks of the two matrices, $r$ and $s$, are both small, the Hadamard product can effectively raise the upper bound of the matrix rank. Therefore, when augmenting the rank of the output features matrix, the choice of $\phi(.)$ becomes less important; what matters is the use of the Hadamard product. This is consistent with the conclusions we obtained from the ablation experiments in the main text. Regardless of what $\phi(.)$ is, the presence of $\phi(.)$ will always enhance the rank of the matrix.
Figure 9: Overall architecture of RAVLT.
Appendix C Implementation Details
Overall Architectures.
Similar to previous works [19, 35, 18], RAVLT employs a four-stage architectural design, as shown in Fig. 9.
Image Classification.
We adopt the training strategy proposed in DeiT [55] with the only supervision is classification loss. Specifically, all models are trained from scratch for 300 epochs with the input resolution of $224\times 224$. Adam is used with a cosine decay learning rate scheduler and 5 epochs of linear warm-up. The initial learning rate, weight decay, and batch-size are set to 0.001, 0.05, and 1024, respectively. We apply the same data augmentation and regularization as DeiT [55] (RandAugment [8] (randm9-mstd0.5-inc1) , Mixup [72] (prob = 0.8), CutMix [71] (prob = 1.0), Random Erasing (prob = 0.25)).
Object Detection and Instance Segmentation.
We apply RetinaNet [42], Mask-RCNN [29], and Cascaded Mask R-CNN [3] as the frameworks based on the MMDetection [5] to evaluate our models. The models are trained under “1 $\times$” (12 training epochs) and “3 $\times$ +MS” (36 training epochs with multi-scale training) settings. For the “1 $\times$” setting, images are resized to the shorter side of 800 pixels while the longer side is within 1333 pixels. For the “3 $\times$ +MS”, multi-scale training strategy is applied to randomly resize the shorter side between 480 to 800 pixels. We use the initial learning rate of 1e-4. For RetinaNet, we set the weight decay to 1e-4. For Mask-RCNN and Cascaded Mask R-CNN, we set it to 5e-2.
Semantic Segmentation.
we implement UperNet [64] and SemanticFPN [36] based on MMSegmentation [7] to validate the models. For UperNet, we follow the previous setting [45] and train the model for 160k iterations with the input size of $512\times 512$. For SemanticFPN, we also use the input resolution of $512\times 512$ but train the models for 80k iterations.